OpenStudy (christos):

Determine which of the following equations is not a function of x. A) x^2 - y = 6 B) y= 3x/x^2-4 C) x^2 - y^2 = 1 D) Square Root of 1-x^3 = y When it says is not a "Function" what does it mean precisely? When a number is a Function of an equation?

OpenStudy (christos):

I might be confuses because the terms are in english which is not my mother language and such

OpenStudy (skullpatrol):

Define "function"

OpenStudy (christos):

An activity or purpose natural to or intended for a person or thing.

OpenStudy (christos):

But I still dont get it :D

OpenStudy (anonymous):

Where are the equations?

OpenStudy (christos):

I will type them wait

hartnn (hartnn):

if you have 1 value of x for 2 values of y, then your relation is not a function. example : x=y^2

hartnn (hartnn):

*2 or more

OpenStudy (anonymous):

its simply tells you to choose the right equation that satisfy x. or vice versa.

OpenStudy (christos):

A) x^2 - y = 6 B) y= 3x/x^2-4 C) x^2 - y^2 = 1 D) Square Root of 1-x^3 = y

OpenStudy (anonymous):

wheres the given?

OpenStudy (christos):

"Determine which of the following equations is not a function of x." Thats the question

OpenStudy (skullpatrol):

Dictionary definition: An activity or purpose natural to or intended for a person or thing. You need a Math dictionary definition

OpenStudy (skullpatrol):

Use your textbook or notes.

OpenStudy (christos):

I see where the fuction is getting at. F(x) and such

OpenStudy (christos):

If I have found the correct explanation that is

OpenStudy (anonymous):

when you put one value in independent variable and you get exactly one corresponding value of dependent variable that equation is called function otherwise not a function. in all the above given equations the equation that is not a function is (C) the rest are functions by definition.

OpenStudy (christos):

hmm

OpenStudy (anonymous):

i get it, first, isolate y. then substitute any number for x. lets say substitute 0 for comfort. then if the equation gives you two values. then, the equation is not a function of x.in your case. letter D is the correct answer(not a function of x). because it gives you -1 and +1 value. as for letter C. it gives you a imaginary number so its negletable.

OpenStudy (skullpatrol):

http://en.wikipedia.org/wiki/Function_(mathematics)

OpenStudy (christos):

@rizwan_uet so how can you determine whenever it's possible to work with dependent/independent variables in one equation? May I assume that you "Can't" Work like that in C) x^2 - y^2 = 1 because both variables have the power of 2?

OpenStudy (christos):

or generally both powers have a power greater than 1 ?

OpenStudy (christos):

Or are there any formulas that I can use to find out?

OpenStudy (christos):

Except if there is a simplier way ofcourse. The simplier the better

OpenStudy (anonymous):

heres my reference. http://answers.yahoo.com/question/index?qid=20080813053127AAPsnFy no theres no formula, just isolate y. use your algebra :)

OpenStudy (christos):

So Long Story Sort: Answer is C because Y is in the power of 2 Correct?

OpenStudy (christos):

Short*